# Humans Hallucinate Too: Language Models Identify and Correct Subjective Annotation Errors With Label-in-a-Haystack Prompts

### v1

@@ 1 Introduction
Figure 1: Label-in-a-Haystack Rectification (LiaHR): The query also appears in the prompt as a demo. The LLM is instructed to perform the actual task, as captured by the label names. We leverage the failure to correctly copy-paste the query’s label for filtering and correction.
@@ 2.2 Error Detection
Previously, error detection has been carried out in a variety of ways and levels of intervention. For example, researchers used the EM algorithm Dawid and Skene (1979) to assign confidence to each annotator’s evaluations Hovy et al. (2013) given an annotated dataset, partly in order to disregard unreliable annotators. However, this risks marginalizing idiosyncratic viewpoints, which may otherwise be internally consistent Chochlakis et al. (2025). Similarly as as a post-processing step, previous work has used trained models on the dataset to assess the quality of the labels, either directly, e.g., with dataset cartography Swayamdipta et al. (2020); Mokhberian et al. (2022), where each data point is mapped onto a 2D space depending on the confidence and the accuracy of the predictions, or indirectly, e.g., with self distillation Stanton et al. (2021). Label verification has also been explored online by using predictions from a model, such as a Large Language Model (LLM), and checking them against the annotations Feng and Narayanan (2024). However, this method trivially considers differing perspectives invalid. Previous work has also shown how the prior biases of LLMs ossify their posterior predictions Chochlakis et al. (2024), which in turn leads to failures in accommodating different perspectives during regular inference. This further narrows the breadth of subjective assessment we ideally want to capture and limits our ability to potentially elicit different predictions from LLMs in subjective settings. Finally, when iterating in batches, verification checks cannot be automated similar to the aforementioned post-processing step due to the lack of sufficient data, so checks need to be manual, such as analyzing disagreement or having annotators engage in consensus talks Paletz et al. (2023), significantly increasing costs. Liu et al. (2023) showed that LLMs do not model ambiguity, an important component of disagreement.
@@ 3 Methodology
First, it is important to provide some working definition of reasonableness (and in turn, what a subjective task is). For our purposes, we consider a document-label pair to be reasonable if and only if a person who would annotate differently can nonetheless consider some reasoning process that leads to that label valid. That is, if a human can agree that a reasoning process is valid, coherent, and faithful Jacovi and Goldberg (2020) with respect to the label, then that label is deemed reasonable111in the case of initial disagreement with a specific rationale, iterative refinement until agreement is achieved is valid, assuming the reasoning remains faithful to the labels.
@@
As shown in Figure 1, the query and its candidate label are included in the prompt as the first example, and the model is instructed to perform the task described by its labels. However, due to inclusion of the label in the prompt already, we essentially check whether the model is able to copy-paste the query’s label onto its output. Given previous results about the collapse of the posterior predictions of LLMs to the priors in complex subjective tasks Chochlakis et al. (2024), we expect that in cases where the gold labels are “judged” to be erroneous by the model, the copy-pasting will fail, flagging a label for further review. In addition to this ability, this setting also allows us to immediately get alternative labels for the example, a property that the baseline does not possess. In this manner, we do not waste data by discarding examples that are flagged by the model.
@@ 3.1 Proxy Properties
In this section, we define and present desirable proxy properties that can be used as proxies for the label filtering and correction ability:
@@
This is the first requirement. Although “small” is nebulous, the model should be copying the gold labels significantly better compared to its performance as a classifier. Having a smaller gap to the dataset’s labels indicates an ability to agree with different perspectives, and it assumes that most of the dataset has been annotated properly.
@@
Respecting different opinions is also an integral property. Here, we also assume that most annotators have annotated most of the dataset properly. For this quality of the model, we can use annotations from different individuals and expect the model to predict reasonableness or successfully copy the labels at equally high rates for them all.
@@ 3.4 Metrics
Because the LiaHR format is identical to classification, we use metrics appropriate for classification to evaluate the performance of copy-pasting and get a more nuanced picture of the predictions of the model. We use Jaccard Score and Micro F1 for multilabel, and accuracy and F1 for single-label cases.
@@ 4.1 Datasets
SemEval 2018 Task 1 E-c Mohammad et al. (2018)
A multilabel emotion recognition benchmark containing annotations for 11 emotions: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, and trust. We use only the English subset. We refer to this as SemEval.
@@
Single-label binary harm dataset, which contains various templates populated with reclaimed LGTBQ+ slurs. It contains two harm labels: assuming in-group and out-group authors. Using one or the other without explicit mention, we can evaluate the Diversity property with a known and controllable causal factor. Moreover, it is challenging because it includes a realistic confounding factor: the interplay between politeness guardrails and our desired behavior. We create splits to be as balanced as possible, but still present ROC-AUC to avoid any bias. Because the labels are binary, we use the opposite label instead of randomizing.
@@ 4.2 Implementation Details
We use the 4-bit quantized versions of the open-source LLMs through the HuggingFace Wolf et al. (2020) interface for PyTorch. We use GPT-3.5 Turbo (gpt-3.5-turbo), GPT-4 (gpt-4-turbo), and GPT-4o (gpt-4o-mini), Llama-2 7B and 70B (meta-llama/Llama-2-#b-chat-hf), and Llama-3 8B and 70B (meta-llama/Meta-Llama-3-#B-Instruct). We chose only finetuned models Ouyang et al. (2022) to avoid confounding factors. We use random retrieval of examples. We train Demux Chochlakis et al. (2023) as the smaller model for ecological validity. When sampling random labels, we ensure at least one label is present (i.e., we do not allow Nones because of their higher plausibility). Results for proxy properties are 3 different seeds with 100 inference examples each, whereas the entire corpus is used for training and evaluation of smaller models. For more details, see Appendix A and B.
@@ 4.3 Evaluating Proxy Properties
We present our results for all222some API-based models were deprecated during the course of our experiments, so we skip them where they are not available. For additional results, such as GPT-4, see Appendix C. models in Figure 2 for LiaHR and Figure 3 for the baseline. In Figure 2, we present the performance of the model on the copy-paste task when using gold (Query w/ gold) and random (Query w/ rand) labels for the demo query, as well as the performance of the model on the gold labels when the query label is random (and therefore the model has not seen the test label for the query; Gold perf w/ rand). All results are shown for 5, 15, 25, 55, and 75 shot (to demonstrate scalability). For Figure 3, we show the first two scenarios, where the document is presented to the LLM with its paired label (Gold pair) or a random label (Rand pair).
In LiaHR, we see clear evidence for our desired behavior in bigger and more capable models, specifically GPT-3.5, GPT-4o, and Llama-3 70b. These models seems to display all the properties we check for: Nonconformity, Rectification, and Noise rejection. First, the Jaccard Score with gold labels for the query is not perfect, yet it is significantly higher to the same model’s performance on the benchmark Chochlakis et al. (2024) (Nonconformity). Then, when we use random labels instead of gold for the query in the prompt, we see performance drop dramatically (Noise rejection). Moreover, it is interesting to see that the predictions more closely match the gold labels than the random labels, even though these random labels are provided in the prompt (Rectification). Additionally, Llama-3 8b also seems to meet some of the properties, though only for a fraction of cases.
@@
Figure 6: Success rate of copying the labels of LiaHR on GoEmotions and MFRC with aggregate labels, random labels, and annotator labels, shown for 15-shot prompts.
@@
In Appendix D, we also show our results and very interesting findings for these three properties in MFRC, where smaller models seem to be treating the gold and random labels similarly.
@@
We examine Diversity separately, in Figures 6 and 7. Figure 6 shows the success rates of copy-pasting on MFRC and GoEmotions between the gold, random, and individual annotator labels, using otherwise the same exact prompts and only differing the labels to avoid confounding factors. We see that, first, all annotators tend to be clustered together with small rejection rates, indicating that the model tends to accept all different perspectives equally. Second, we see that their performance is better compared to random. Finally, the similarity between the annotators shown can be very low (e.g., as low as 0.433 Jaccard Score on GoEmotions between annotators), representing consistently different perspectives. All these pieces of evidence indicate that most models achieve Diversity. Moreover, we see a marked difference between annotators and the aggregate, with the latter displaying higher rejection rates, indicating that part of our aforementioned results on MFRC and GoEmotions can be explained as aggregation artifacts Chochlakis et al. (2025).
Figure 7 shows that LiaHR can successfully accept both in-group and out-group perspectives in the QueerReclaimLex benchmark without explicit prompting, instead learning implicit causal cues from few examples. Results show that the models tend to model out-group annotations better. However, more capable models also recognize reclaimed slurs as not harmful when used by in-group speakers, scaling performance with more demonstration, indicating the robustness of LiaHR to the guardrails placed on models.
4.4 Human Evaluation
Results for our human evaluations are presented in Table 1 for SemEval for the models that meet our defined properties. More detailed results on SemEval and GoEmotions can be found in Appendix B. We see that Llama-3 70b and GPT-3.5 do not show enough discriminability between reasonable and unreasonable labels, although their results are strong in terms of preference for their labels when the copy-paste task was performed incorrectly. However, GPT-4 and 4o can distinguish between reasonable and unreasonable labels and also propose better alternatives for unreasonable labels. The results show strong statistical significance, but also large effect sizes. This is not the case when checking for the ICL prediction of the models. This shows that the predictions of LLMs are not preferred over the gold labels by humans, indicating that our settings are important to achieve proper filtering. We also see that the explicit baseline shows sufficient discriminability for both Llama-3 70b and GPT-4o.
4.5 Ecological Validity
In addition to the human evaluations and defining and evaluating proxy properties, we also perform ecological validity studies, and compare to other online methods. That is, even though we have shown the models have desirable properties, and people tend to prefer them over the original labels, do models trained on them perhaps show erratic behavior? For all the settings introduced in Section 3.3, we show the results in Table 2 (additional results in Appendix G). The results indicate that the new labels lead to slightly better generalization performance, although the methods need to be applied throughout the annotation process to get the maximum benefit. Note that SemEval is a smaller dataset, leading to extra performance decreases when examples are filtered instead of corrected. Noticeably, we also see that using the raw predictions of the models leads to substantial deterioration in performance. In addition to the humans evaluations, these results indicate that our proposal for “reasonableness” checks rather than simply using the LLM as classifier is warranted.
@@ 4.5 Ecological Validity
[…]
@@ 5 Conclusion
In this work, we propose “reasonableness” checks to improve the signal-to-noise ratio in subjective language annotations. We leverage LLMs and introduce LiaHR, which is able to both filter and correct unreasonable annotations, and a simple baseline that detects unreasonable annotations. We demonstrate that both approaches satisfy desirable proxy properties, pass human evaluations, and show ecological validity when used to train smaller models. Moreover, we show that the model can pick up on causal yet implicit cues from few examples reliably. To further corroborate our findings on LiaHR, we also show how it performs in objective tasks in Appendix E, an analysis of the copy-paste performance across shots, model families and sizes in Appendix F, that individual labels are uniformly affected in Appendix H, and the robustness to the position of the query in Appendix I.
6 Limitations
We want to emphasize that our model is not an oracle. The model does not provide ground truth / gold labels and could be biased in other ways. While our experiments show that humans prefer the model’s labels when it is performing correction, we advocate for additional checks when LiaHR is used during the annotation process. For example, if some submitted labels for a specific example do not pass the LiaHR filter, instead of always using its alternative predictions, the same document can be shown to the annotator at a later stage to verify and potentially correct the label themselves.
Secondly, we want to note that, despite the remarkable robustness of the framework on the reclaim slurs dataset, QueerReclaimLex, its performance on the in-group data is noticeably worse than the out-group. This indicates that there might be some bias in the decisions of the model. Therefore, we urge immense caution when the framework is used in sensitive settings.

### v2

@@ 1 Introduction
Figure 1: Label-in-a-Haystack Rectification (LiaHR): The query also appears in the prompt as a demo. The LLM is instructed to perform the actual task, as captured by the label names. We leverage the failure to correctly copy-paste the query’s label to flag the query-label pair, for filtering or even correction based on the prediction.
@@ 2.2 Error Detection
Previously, error detection has been carried out in a variety of ways and levels of intervention. One research thread assumes a single correct answer per item, and proceed to identify errors or “spammer” annotators. Examples include the Dawid-Skene algorithm Dawid and Skene (1979), MACE Hovy et al. (2013), and CrowdKit Ustalov et al. (2021). However, these methods fail the basic assumption of our work, as they do allow difference in opinion, marginalizing idiosyncratic viewpoints, which may otherwise be internally consistent Chochlakis et al. (2025). Similar approaches that allow for disagreement still assign scores per item and annotator individually and not for separately for each pair, like CrowdTruth Dumitrache et al. (2018).
In another research thread, again as a post-processing step, previous work has used trained models on the dataset to assess the quality of the labels, either directly, e.g., with dataset cartography Swayamdipta et al. (2020); Mokhberian et al. (2022), where each data point is mapped onto a 2D space depending on the confidence and the accuracy of the predictions, or indirectly, e.g., with self distillation Stanton et al. (2021).
Label verification has also been explored online by using predictions from a model, such as a Large Language Model (LLM), and checking them against the annotations Feng and Narayanan (2024). However, this method trivially considers differing perspectives invalid. Previous work has also shown how the prior biases of LLMs ossify their posterior predictions Chochlakis et al. (2024), which in turn leads to failures in accommodating different perspectives during regular inference. This further narrows the breadth of subjective assessment we ideally want to capture and limits our ability to potentially elicit different predictions from LLMs in subjective settings. When iterating in batches, verification checks cannot be automated similar to the aforementioned post-processing step due to the lack of sufficient data, so checks need to be manual, such as analyzing disagreement or having annotators engage in consensus talks Paletz et al. (2023), significantly increasing costs. Liu et al. (2023) showed that LLMs do not model ambiguity, an important component of disagreement.
@@ 3 Methodology
First, it is important to provide some working definition of reasonableness (and in turn, what a subjective task is). For our purposes, we consider a document-label pair to be reasonable if and only if a person who would annotate differently can nonetheless consider some reasoning process that leads to that label valid. That is, if a human can agree that a reasoning process is valid, coherent, and faithful Jacovi and Goldberg (2020) with respect to the label, then that label is deemed reasonable111in the case of initial disagreement with a specific rationale, iterative refinement until agreement is achieved is valid, assuming the reasoning remains faithful to the labels. We present the general and intuitive description of our methods in this section and a more mathematically rigorous description in Appendix A.
@@
As shown in Figure 1, the query and its candidate label are included in the prompt as the first example, and the model is instructed to perform the task described by its labels. However, due to inclusion of the label in the prompt already, we essentially check whether the model is able to copy-paste the query’s label onto its output. Given previous results about the collapse of the posterior predictions of LLMs to the priors in complex subjective tasks Chochlakis et al. (2024), we expect that in cases where the gold labels are “judged” to be erroneous by the model, the copy-pasting will fail, flagging a label for further review. In addition to this ability, this setting also allows us to immediately get alternative labels for the example, a property that the baseline does not possess. In this manner, we do not waste data by discarding examples that are flagged by the model. We note that when using random labels for the query document, we sample them from a random document in the dataset, similar to the baseline.
@@ 3.1 Proxy Properties
In this section, we define and present desirable proxy properties that can be used as proxies for the label filtering and correction ability that practitioners can use to guide model selection. Note that since strictness is not required because of their proximate nature, some of them are fuzzy and heuristic.
@@
This is the first requirement. Although “small” is nebulous, the model should be copying the gold labels significantly better compared to its performance as a classifier. Having a smaller gap to the dataset’s labels indicates an ability to agree with different perspectives, and it assumes that most of the dataset has been annotated properly222in this assumption, we take for granted that annotators have been screened, trained, attention-checked, etc. Namely, we assume quality data collection.
@@
Respecting different opinions is also an integral property. Here, we also assume that most annotators have annotated most of the dataset properly333we make the same assumptions as above. For this quality of the model, we can use annotations from different individuals and expect the model to predict reasonableness or successfully copy the labels at equally high rates for them all.
@@ 3.4 Metrics
Because the LiaHR format is identical to classification, we use classification metrics to evaluate the performance of copy-pasting and get a more nuanced picture of the predictions of the model. We use Jaccard Score and Micro F1 for multilabel, and accuracy and F1 for single-label cases.
@@ 4.1 Datasets
SemEval 2018 Task 1 E-c (SemEval; Mohammad et al. 2018)
A multilabel emotion recognition benchmark containing annotations for 11 emotions: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, and trust. We use only the English subset.
@@
Single-label binary harm dataset, which contains various templates populated with reclaimed LGTBQ+ slurs. It contains two harm labels: assuming in-group and out-group authors. Using one or the other without explicit mention, we can evaluate the Diversity property with a known and controllable causal factor. This setting serves as a stress test, since reclaimed slurs in general are a documented failure case for, e.g., toxicity classifiers Sap et al. (2019); Haimson et al. (2021); Sap et al. (2022), allowing us to examine whether systematic biases in LLMs influence their decisions in our framework. For the same reasons, it is challenging because it includes a realistic confounding factor: the interplay between politeness guardrails and our desired behavior, as slurs are explicitly included throughout the prompt. We create splits to be as balanced as possible, but also present ROC-AUC to avoid bias. Because the labels are binary, we use the opposite label instead of randomizing the query label.
@@ 4.2 Implementation Details
We use the 4-bit quantized versions of the open-source LLMs through the HuggingFace Wolf et al. (2020) and bitandbytes interface for PyTorch. We use GPT-3.5 Turbo (gpt-3.5-turbo), GPT-4 (gpt-4-turbo), and GPT-4o (gpt-4o-mini), Llama-2 7B and 70B (meta-llama/Llama-2-#b-chat-hf), and Llama-3 8B and 70B (meta-llama/Meta-Llama-3-#B-Instruct). We chose only finetuned models Ouyang et al. (2022) to avoid confounding factors. We use random retrieval of examples. We train Demux Chochlakis et al. (2023) as the smaller model for ecological validity. When sampling random labels, we ensure at least one label is present (i.e., we do not allow Nones because of their higher plausibility). Results for proxy properties are 3 different seeds with 100 inference examples each. The entire corpus is used for training and evaluation of smaller models. Unless otherwise noted, we show 95% confidence interval around the mean. For more details, see Appendix B and C.
@@ 4.3 Evaluating Proxy Properties
We present our results for all444some API-based models were deprecated during the course of our experiments, so we skip them where they are not available. For additional results, such as GPT-4, see Appendix D. models in Figure 2 for LiaHR and Figure 3 for the baseline. In Figure 2, we present the performance of the model on the copy-paste task when using gold (Query w/ gold) and random (Query w/ rand) labels for the demo query, as well as the performance of the model on the gold labels when the query label is random (and therefore the model has not seen the test label for the query; Gold perf w/ rand). All results are shown for 5, 15, 25, 55, and 75 shot (to demonstrate scalability). For Figure 3, we show the first two scenarios, where the document is presented to the LLM with its paired label (Gold pair) or a random label (Rand pair).
In LiaHR, we see clear evidence for our desired behavior in bigger and more capable models, specifically GPT-3.5, GPT-4o, and Llama-3 70b. These models seem to display all the properties we check for: Nonconformity, Rectification, and Noise rejection. First, the success rate with gold labels for the query is not perfect (meaning 1.0), yet it is significantly higher compared to the same model’s performance on the benchmark Chochlakis et al. (2024). This means that the model does not conform to the gold labels completely, yet is greatly influenced by them in its predictions (otherwise we would anticipate performance much closer to its “regular” predictions). By meeting both these criteria, the aforementioned models meet the Nonconformity property. Then, when we use random labels instead of gold for the query in the prompt, we see the success rate drop dramatically compared to when gold labels are presented (that it, when comparing Query w/ gold to Query w/ rand). This indicates that models achieve the Noise Rejection property. Moreover, it is interesting to see that, when random labels are provided, the predictions match more closely the gold labels (Gold perf w/ rand) than these random labels (Query w/ rand). Since this criterion is met, the models achieve Rectification.
@@
Figure 6: Success rate of copying the labels of LiaHR on GoEmotions and MFRC with aggregate labels, random labels, and annotator labels (Ann#), shown for 15-shot prompts. For GoEmotions, actual annotator IDs are: Ann0 = 37, Ann1 = 4, Ann2 = 61. For MFRC: Ann0 = 0, Ann1 = 1, Ann2 = 3.
@@
In Appendix E, we also show our results and very interesting findings for these three properties in MFRC, where smaller models seem to be treating the gold and random labels similarly.
BERT-based baseline
In Appendix L, we show results for a BERT-based filtering baseline, showing it underperforms 5-shot GPT-4o while requiring the entire dataset to be trained, disincentivizing its usage from beginning to end of the data collection.
@@
We examine Diversity separately, in Figures 6 and 7. Figure 6 shows the success rates of copy-pasting on MFRC and GoEmotions between the gold, random, and individual annotator labels, using otherwise the same exact prompts and only differing the labels to avoid confounding factors. We first see that all annotators tend to be clustered together with small rejection rates, indicating that the model tends to accept all different perspectives equally. Second, we see that their performance is better compared to random. Finally, the similarity between the annotators shown can be very low (e.g., as low as 0.433 Jaccard Score on GoEmotions between annotators), representing consistently different perspectives. Consequently, the majority of the disagreement between annotators is being preserved by the model organically, without any intervention. All these pieces of evidence indicate that most models achieve Diversity. Moreover, we see a marked difference between annotators and the aggregate, with the latter displaying higher rejection rates, indicating that part of our aforementioned results on MFRC and GoEmotions can be explained as aggregation artifacts Chochlakis et al. (2025).
Figure 7 shows that LiaHR can successfully accept both in-group and out-group perspectives in the QueerReclaimLex benchmark without explicit prompting, instead learning implicit causal cues from few examples. Results show that the models tend to model out-group annotations better. However, more capable models also recognize reclaimed slurs as not harmful when used by in-group speakers, scaling performance with more demonstration, indicating the robustness of LiaHR to the guardrails placed on models, and an ability to counterbalance systematic biases with few demonstrations, a challenging problem in toxicity classifiers with reclaimed slurs Sap et al. (2019); Haimson et al. (2021); Sap et al. (2022). Thus, LiaHR proves robust to our stress test for whether the model itself might introduce biases in the dataset.
@@
Original
$0.652{\scriptstyle\pm 0.001}$
$0.689{\scriptstyle\pm 0.002}$
Replaced
$\textbf{0.653}{\scriptstyle\pm 0.000}$
$\textbf{0.692}{\scriptstyle\pm 0.003}$
Replaced (trn)
$0.642{\scriptstyle\pm 0.001}$
$0.680{\scriptstyle\pm 0.002}$
Filtered
$0.652{\scriptstyle\pm 0.002}$
$0.679{\scriptstyle\pm 0.002}$
Bsl Filtered
$0.638{\scriptstyle\pm 0.001}$
$0.680{\scriptstyle\pm 0.003}$
Predictions
$0.427{\scriptstyle\pm 0.002}$
$0.613{\scriptstyle\pm 0.000}$
@@
4.4 Human Evaluation
Results for our human evaluations are presented in Table 1 for SemEval for the models that meet our defined properties. More detailed results on SemEval and GoEmotions can be found in Appendix C. We see that Llama-3 70b and GPT-3.5 do not show enough discriminability between reasonable and unreasonable labels, although their results are strong in terms of preference for their labels when the copy-paste task was performed incorrectly. However, GPT-4 and 4o can distinguish between reasonable and unreasonable labels and also propose better alternatives for unreasonable labels. The results show strong statistical significance, but also large effect sizes. This is not the case when checking for the ICL prediction of the models. This shows that the predictions of LLMs are not preferred over the gold labels by humans, indicating that our settings are important to achieve proper filtering. We also see that the explicit baseline shows sufficient discriminability for both Llama-3 70b and GPT-4o.
4.5 Ecological Validity
In addition to the human evaluations and defining and evaluating proxy properties, we also perform ecological validity studies, and compare to other online methods. That is, even though we have shown the models have desirable properties, and people tend to prefer them over the original labels, do models trained on them perhaps show erratic behavior? For all the settings introduced in Section 3.3, we show the results in Table 2 (additional results in Appendix H). The results indicate that the new labels lead to slightly better generalization performance, although the methods need to be applied throughout the annotation process to get the maximum benefit. Note that SemEval is a smaller dataset, leading to extra performance decreases when examples are filtered instead of corrected. Noticeably, we also see that using the raw predictions of the models leads to substantial deterioration in performance. In addition to the humans evaluations, these results indicate that our proposal for “reasonableness” checks rather than simply using the LLM as classifier is warranted.
@@ 5 Conclusion
In this work, we propose “reasonableness” checks to improve the signal-to-noise ratio in subjective language annotations. We leverage LLMs and introduce LiaHR, which is able to both filter and correct unreasonable annotations, and a simple baseline that detects unreasonable annotations. We demonstrate that both approaches satisfy desirable proxy properties, pass human evaluations, and show ecological validity when used to train smaller models. Moreover, we show that the model can pick up on causal yet implicit cues from few examples reliably. While our experiments show that humans prefer the model’s labels when it is performing correction, we advocate for usage during the annotation process, with additional checks. For example, if some submitted labels for a specific example do not pass the LiaHR filter, instead of always using its alternative predictions, the same document can be shown to the annotator at a later stage to verify and potentially correct the label themselves.
To further corroborate our findings on LiaHR, we also show how it performs in objective tasks in Appendix F, an analysis of the copy-paste performance across shots, model families and sizes in Appendix G, that individual labels are uniformly affected in Appendix I, and the robustness to the position of the query in Appendix J.
6 Limitations and Ethics
We want to emphasize that our model is not an oracle. The model does not provide ground truth / gold labels and could be biased in other ways.
Our work entails some potential for deliberate misuse. Although we advocate for using individual perspectives as demonstrations in LiaHR throughout our work, deliberate misuse might include skewing the perspectives in the prompt and using the rejection from LiaHR as justification for rejecting minority labels and preventing certain valid perspectives from entering the data (i.e., gate-keeping). Therefore, we want to emphasize that LiaHR assessments can only be considered valid (though not necessarily correct) when the perspective being evaluated (the query label) coincides with the perspective in the demonstrations. The predictions of the model should not be taken into account otherwise.
Accidental misuse includes model biases seeping into the labels. We want to note that, despite the remarkable robustness of the framework on the reclaim slurs dataset, QueerReclaimLex, its performance on the in-group data is noticeably worse than the out-group. This indicates that there might be some bias in the decisions of the model. Moreover, assessments of harm are inherently subjective, reflecting differences in individual and cultural perceptions. The original work aggregates distinct gender identities (e.g., non-binary, transgender) under the umbrella term gender-queer and treats them as largely synonymous. While this simplification overlooks the diversity of perspectives, we follow the original work’s adoption of this grouping as a pragmatic choice to support our analysis. As understandings of differing perspectives continue to evolve, future work should aim to incorporate a broader pool of annotators and explore methods for capturing the nuance and variability of perceived harm across different communities. Therefore, we urge immense caution when the framework is used in sensitive settings.